Simplify the following expressions: (a) (cos π/4 + i sin π/4)(cos 3π/4 + i sin 3π/4), (b)
√ 3, (b) w = −2 1 √
(cos π/4 + i sin π/4)2 (cos π/6 + i sin π/6)
1
Expert's answer
2020-05-28T17:55:24-0400

a) "(cos (\u03c0\/4) + i sin (\u03c0\/4))(cos (3\u03c0\/4) + i sin (3\u03c0\/4))="

"=(cos (\u03c0\/4) + i sin (\u03c0\/4))(-cos (\u03c0\/4) + i sin (\u03c0\/4))="

"=-sin^2(\pi/4)-cos^2(\pi/4)=-1"


b) "(cos (\u03c0\/4) + i sin (\u03c0\/4))^2 (cos (\u03c0\/6) + i sin (\u03c0\/6))="

"=(cos (2\u03c0\/4) + i sin (2\u03c0\/4)) (cos (\u03c0\/6) + i sin (\u03c0\/6))="

"= i (cos (\u03c0\/6) + i sin (\u03c0\/6))=-sin (\u03c0\/6)+icos (\u03c0\/6)="

"=-1/2+i\sqrt{3}/2"